OpenStudy (anonymous):

ANSWER IN 10 MINUTES FOR A MEDAL + FAN!! ANSWER IN 10 MINUTES FOR A MEDAL + FAN!! Which is an equation in point-slope form for the given point and slope? Point: (8, -4) ; Slope: -5/6

OpenStudy (anonymous):

I WILL MEDAL / FAN!!!

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = -5/6 (8,-4)...x1 = 8 and y1 = -4 now we sub y - (-4) = -5/6(x - 8) y + 4 = -5/6(x - 8)